# Fault-Tolerant Decentralized Control for Large-Scale Inverter-Based Resources for Active Power Tracking

### v1

@@
Abstract
@@ 1 Introduction
The paper is organized as follows: Section-2 introduces the mathematical notations used throughout the paper and the preliminaries to better understand the control development. In Section-3, the inverter model is presented followed by the low-level and high-level control design in Section-4. Section-5 presents the simulation results of the designed controller and the results are compared with the results in the literature.
@@ 4.1 Low-Level Control
It follows that $\tilde{\mathbf{i}}(t)\in\mathbb{L}_{2}$. From the implications $\tilde{\mathbf{i}}(t)$ is uniformly continuous. Thus, invoking Barbalat’s lemma [15] it follows that $\tilde{\mathbf{i}}(t)\longrightarrow\underline{\mathbf{0}}$.
Proposition 4.3.
@@
Thus, $\mathbf{i}(t)\in\mathcal{B}\bigg{(}\mathbf{i}_{m}(t),\left\lVert\tilde{\mathbf{i}}(0)\right\rVert+\frac{\overline{r}}{\sqrt{2\gamma_{r}}}\bigg{)}$.
From the Proposition 4.3, the resulting power from the single IBR satisfies
@@ 6 Conclusion and Future Work
Acknowledgment
This work was supported by the U.S. Department of Energy (DOE) under Award DE-EE0009340. This article reflects only the author’s views and DOE is not responsible for any use that may be made of the information it contains.
Conflict of Interest
Authors have no conflict of interest to declare.
Data Availability Statement
The datasets generated and supporting the findings of this article are obtainable from the corresponding author upon reasonable request.

### v2

@@
Fault-Tolerant Decentralized Control for Large-Scale Inverter-Based Resources for Active Power Tracking
Satish Vedula, Ayobami Olajube, and
Olugbenga Moses Anubi
Department of Electrical and Computer Engineering, the Center for Advanced Power Systems, Florida State University
E-mail: {svedula, aolajube, oanubi}@fsu.edu
1 Abstract
@@ 2 Introduction
The paper is organized as follows: Section-3 introduces the mathematical notations used throughout the paper and the preliminaries to better understand the control development. In Section-4, the inverter model is presented followed by the low-level and high-level control design in Section-5. Section-6 presents the simulation results of the designed controller and the results are compared with the results in the literature.
@@ 5.1 Low-Level Control
It follows that $\tilde{\mathbf{i}}(t)\in\mathbb{L}_{2}$. From the implications $\tilde{\mathbf{i}}(t)$ is uniformly continuous. Thus, invoking Barbalat’s lemma [15] it follows that $\tilde{\mathbf{i}}(t)\longrightarrow\underline{\mathbf{0}}$. ∎
Proposition 1.
@@
Thus, $\mathbf{i}(t)\in\mathcal{B}\bigg{(}\mathbf{i}_{m}(t),\left\lVert\tilde{\mathbf{i}}(0)\right\rVert+\frac{\overline{r}}{\sqrt{2\gamma_{r}}}\bigg{)}$. ∎
From the Proposition 1, the resulting power from the single IBR satisfies